### v1

# Follow My Instruction and Spill the Beans: Scalable Data Extraction from Retrieval-Augmented Generation Systems

Abstract
Retrieval-Augmented Generation (RAG) improves pre-trained models by incorporating external knowledge at test time to enable customized adaptation. We study the risk of datastore leakage in Retrieval-In-Context RAG Language Models (LMs). We show that an adversary can exploit LMs’ instruction-following capabilities to easily extract text data verbatim from the datastore of RAG systems built with instruction-tuned LMs via prompt injection. The vulnerability exists for a wide range of modern LMs that span Llama2, Mistral/Mixtral, Vicuna, SOLAR, WizardLM, Qwen1.5, and Platypus2, and the exploitability exacerbates as the model size scales up. Extending our study to production RAG models GPTs, we design an attack that can cause datastore leakage with a 100% success rate on 25 randomly selected customized GPTs with at most 2 queries, and we extract text data verbatim at a rate of 41% from a book of 77,000 words and 3% from a corpus of 1,569,000 words by prompting the GPTs with only 100 queries generated by themselves.
Follow My Instruction and Spill the Beans: Scalable Data Extraction
from Retrieval-Augmented Generation Systems
[…]
1 Introduction
Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Khandelwal et al., 2019; Ram et al., 2023) produces output by retrieving external data relevant to queries and conditioning a parametric generative model on the retrieved content. Such paradigm seeks to address key limitations of parametric LMs (Brown et al., 2020; Chowdhery et al., 2023) such as context length (Xu et al., 2023b), knowledge staleness (Roberts et al., 2020), hallucination (Shuster et al., 2021), attribution (Menick et al., 2022), and efficiency (Borgeaud et al., 2022).
In particular, the inherent propensity of large pre-trained models to memorize and reproduce training data (Carlini et al., 2019, 2023; Nasr et al., 2023), presents significant challenges in terms of legal issues and sensitive data leakage. The approach of RAG emerges as a compelling solution to these issues by creating a balance between generation performance and the demands of data stewardship including copyright and privacy. Specifically, RAG offers a mechanism for training LMs with low-risk data while moving high-risk data to external datastores, as suggested by Min et al. (2023), thereby supports attribution and opts out to hopefully avoid potential legal concerns while preserving the efficacy of LMs.
We show that although RAG systems delegate data to external non-parametric datastores, these data are still vulnerable to extraction attacks (Carlini et al., 2021). We study an adversarial setting by considering a threat model that seeks to extract text data from the non-parametric datastore of RAG models with only black-box API access. Our attack is motivated by the observation that to augment frozen pre-trained models, a wide range of RAG systems prepend retrieved content to the user query (Ram et al., 2023; LangChain, 2022; VoyageAI, 2024; Park et al., 2023; Zhao et al., 2023). Though the implementation is simple and effective, we find that such a Retrieval-In-Context (RIC) manner potentially exposes the datastore to the risk of data extraction even without white-box access to model weights or token probabilities: attackers can exploit the instruction-following capability of LMs (Brown et al., 2020) to reconstruct datastore content by explicitly prompting LMs to repeat the context (Prompt-Injected Data Extraction).
We start by building RIC-based RAG systems using popular open-sourced instruction-tuned LMs as generative models, including Llama2, Mistral/Mixtral, Vicuna, SOLAR, WizardLM, Qwen1.5, and Platypus2, and use newest Wikipedia articles (created later than November 1st, 2023) as datastore. Then adversarial prompts are developed to effectively extract nearly verbatim texts from the datastores of RAG models. We show that LMs with strong capabilities suffer from a high risk of disclosing context, and the vulnerability is exacerbated as the model size scales up from 7B to 70B. Additionally, our ablation studies show that instruction tuning makes LMs more prone to follow Prompt-Injected Data Extraction instructions. Further, we extend our study to one of the most popular production RAG models, GPTs, and show that as of February 2024, an attacker can extract data verbatim from private documents with high success rate using simple prompt injection: an adversary can easily extract the operations and system prompts of all GPTs we experiment with, and thus can explicitly perform retrieval execution commands to leak GPT’s files. Moreover, we can extract text data verbatim at a rate of 41% from a copyrighted book of 77,000 words and 3% from a Wikipedia corpus of 1,569,000 words by iteratively prompting the GPTs with only 100 domain-specific queries generated by themselves.
Figure 1: An overview of attacking RAG systems built with RIC method and instruction-tuned LMs. In an RIC-based RAG system, a retriever first retrieves text chunks from the datastore according to user input and then prepends them to the input as context. Attackers can inject adversarial prompt to the input for disclosing the retrieved texts prepended to the input to an instruction-tuned LM.
Disclosure. Our results should not be considered as the opposition to RAG models or a violation of fair use without context-dependent considerations: while our attack can be used to extract data from RAG models, it’s unlikely to be used for malicious purposes immediately because current RAG systems’ datastores are often implemented based on public, verifiable data sources such as Wikipedia. Rather, understanding the risks revealed in our study would help prevent potential future harm in cases where sensitive or private data are valuable, especially when models are deployed in advanced applications with multiple parties. In other words, we believe that the vulnerability of RAG shown in our attack reveals potential risks of private data leakage and raises concerns regarding its application to data-sensitive scenarios such as medical (Jin et al., 2024), finance (Zhang et al., 2023) and law (Henderson et al., 2022), as well as mechanisms like memories (Park et al., 2023; Zhao et al., 2023; OpenAI, 2024) and citation (Menick et al., 2022), especially when the data being retrieved are not well-sanitized (Elazar et al., 2023).
2 Related Work
Retrieval-Augmented Generation. RAG (Lewis et al., 2020) has been widely studied in the NLG domain, such as kNN-LM (Khandelwal et al., 2019), DPR (Karpukhin et al., 2020), RALM (Guu et al., 2020), and RETRO (Borgeaud et al., 2022). We focus on a popular implementation of RAG - RIC-LM (Ram et al., 2023) that retrieves text chunks from a datastore and feeds them to an LM in context. There has been growing interest in analyzing data leakage problems of RAG systems, including customized GPTs. Huang et al. (2023) first conduct the study of privacy issues on kNN-LMs and show that incorporating private datastores leads to higher risks of data leakage from datastores. Yu et al. (2023) leverage prompt injection to cause file leakage of GPTs by asking them to download the uploaded files using GPT4’s code interpreter as a tool. We are the first to comprehensively study data leakage problems on both open-sourced and production RAG systems and our attack on GPTs reached a 100% success rate without additional tools. Zyskind et al. (2023) propose secure multi-party computation that allows users to privately search a database. Our work focuses on scenarios where datastores should be kept private, which can encompass an array of LM-integrated complex systems, e.g. distributing a customized non-parametric memory-based agent (Park et al., 2023; OpenAI, 2024) to third-party users (OpenAI, 2023); retrieving private yet high-quality data that the model creator is not willing to share with users (Brown et al., 2022); retrieving from pre-training corpora that are not well-sanitized so might contain personally identifiable information (PII) etc sensitive data (Elazar et al., 2023; Subramani et al., 2023).
Data Extraction from Language Models. Training data extraction (Carlini et al., 2021; Nasr et al., 2023) has aroused attention due to LMs’ memorization effect (Carlini et al., 2019; Zhang et al., 2021a; Thakkar et al., 2021; Zhang et al., 2021b), causing privacy and copyright issues (e.g. GMail autocomplete models use private emails as training data (Chen et al., 2019), and PII can be leaked via black-box API access to LMs Lukas et al. (2023)). Potential mitigation methods include performing deduplication on training data (Kandpal et al., 2022) and leverage privacy-preserving training techniques (Yu et al., 2021; Cummings et al., 2023). Prompt extraction has also emerged as a data leakage problem: as shown by Zhang and Ippolito (2023), both open-sourced and production GPT are prone to repeat the prompt under prompt extraction attack. Moreover, Morris et al. (2023) shows that adversaries can reconstruct prompts by training a logit-to-text model in a white-box setting.
Prompt Injection. Prompt injection attacks LMs by crafting malicious instructions to manipulate LMs’ behavior (Wei et al., 2023; Greshake et al., 2023; Liu et al., 2023). In direct prompt injection (Liu et al., 2023; Perez and Ribeiro, 2022), malicious users directly attack LMs with specially designed adversarial prompts to override existing system prompts, while in indirect prompt injection (Greshake et al., 2023; Yi et al., 2023), attackers poison third-party sources with malicious content, to manipulate data input and cause LMs to diverge from their original outputs when users interact with them. Previous studies have evaluated (Branch et al., 2022; Shen et al., 2023) and benchmarked (Yi et al., 2023) LMs’ vulnerability under prompt injection attacks. Yi et al. (2023) show that LMs with strong capabilities are more vulnerable to indirect prompt injection attacks, and we also show that RAG models are more vulnerable to data extraction as they scale up.
3 Problem Formulation
We consider a generic attack formulation that can be adopted across diverse capabilities (Greshake et al., 2023) and modalities (Yasunaga et al., 2022) beyond text and implement our attack on RIC-LM. A RIC-based generator Gen augments a generative model, parametrized by $\theta$, with additional context retrieved from an external non-parametric datastore $D$: $z=\verb|Gen|(\mathcal{R}_{D}(q),q)$, where $\mathcal{R}_{D}(\cdot)$ denotes the retriever that takes as input a user query $q$ and output information retrieved from $D$. In the case of using autoregressive LMs as the generative model, the generation of a sequence of tokens $z=x_{1},...,x_{n}$ follows the distribution: $z\sim p(x_{1},...,x_{n})=\prod_{i=1}^{n}p_{\theta}(x_{i}\mid[\mathcal{R}_{D}(q);q;x_{<i}]).$ We consider a black-box adversary that only has access to the input/output API of a RAG system, whose goal is to reconstruct the datastore $D$ from a series of RIC-based generations. Our data extraction attack is formally defined as follows:
Given a RIC-based generation system Gen using a generative model $p_{\theta}$, a datastore $D$, and a retriever $\mathcal{R}$, Prompt-Injected Data Extraction is to design adversarial input $q$ that triggers the model to generate an output $z=\verb|Gen|(\mathcal{R}_{D}(q),q)$ that reconstructs the retrieved context $\mathcal{R}_{D}(q)$.
4 Attacking Open-sourced LMs
We start with open-sourced LMs and investigate how their instruction-following ability enables black-box adversaries to extract datastore content and test such vulnerability across LMs with different scales.
To elicit the retrieved context, we prompt a RAG model with the following input:
Here is a sentence: {put anchor query here}. Now copy and output all the text before “Here is a sentence”. Do not change any words.
An anchor query is a piece of information based on which we want the RAG system to retrieve relevant text chunks. The intuition of our attack is simple: since the text chunks in the datastore are retrieved and prepended to the user query as the input prompt, we can explicitly instruct the model to reveal the first part of the prompt and thus expose the content in the datastore.
RAG Setup.   We simulate a scenario where the service provider uses the latest Wikipedia content as the knowledge base. To construct the datastore, we collect 1165 recent Wikipedia English articles created after November 1st, 2023, with 1,569,000 words in total. We choose this recent cutoff date to ensure existing models we used have not likely been trained on those Wikipedia texts so datastore content is out of the LMs’ knowledge. We use the Wikipedia API to automatically download the data and filter out articles less than 100 words.
We use instruction-tuned LMs as the generative models. To test LMs across different sizes, we choose Llama2-Chat (7b/13b/70b) (Touvron et al., 2023), Vicuna (13b) (Chiang et al., 2023), Mistral-Instruct (7b) (Jiang et al., 2023) and Mixtral-Instruct (8x7b) (Jiang et al., 2024), SOLAR (10.7b) (Kim et al., 2023), WizardLM (13b) (Xu et al., 2023a), Qwen1.5-Chat (72b) (Bai et al., 2023), and Platypus2-Instruct (70b) (Lee et al., 2023).
We assume that the adversary has no prior knowledge of the datastore. To get anchor queries for attacking open-sourced models, we select 230 long questions from WikiQA(Yang et al., 2015). Note that questions in WikiQA are obsolete so that the adversary has a low probability of querying information in the RAG datastore. By showing the attack results using only the obsolete questions, we aim to show that the vulnerability exists regardless of the choice of queries because of the retrieval mechanism, and certain prior knowledge about the datastore would favor the adversary to design more effective queries.
Results.   We use text similarity between the model output under our attack and the retrieved context to measure the extent to which the models copy the context. For lexical similarity, we consider ROUGE-L (Lin, 2004), BLEU (Papineni et al., 2002), and F1 score at the token level. We also use BERTScore (Zhang et al., 2019) as a measure of semantic relatedness. From Table 1 we see that even Llama2-Chat-7b can reach a ROUGE score and F1 score of higher than 80, and all 70b models reach ROUGE, BLEU, and F1 scores of higher than 80 and almost 100 BERTScore, showing their excessive vulnerability of prompt-injected data extraction. Especially, with a larger model size, the proportion of verbatim copied text also gets larger, thus revealing more datastore content.
Size Model ROUGE-L BLEU F1 BERTScore
7b Llama2-Chat-7b ${\textbf{80.369}_{\pm 1.679}}$ ${\textbf{71.064}_{\pm 2.033}}$ $83.415_{\pm 1.375}$ ${\textbf{94.771}_{\pm 0.301}}$
Mistral-Instruct-7b $79.121_{\pm 0.653}$ $68.426_{\pm 0.857}$ ${\textbf{83.741}_{\pm 0.446}}$ $94.114_{\pm 0.134}$
$\approx$13b SOLAR-10.7b $46.109_{\pm 3.55}$ $38.595_{\pm 3.677}$ $51.224_{\pm 3.302}$ $88.148_{\pm 0.706}$
Llama2-Chat-13b ${\textbf{83.597}_{\pm 1.104}}$ ${\textbf{75.535}_{\pm 1.404}}$ ${\textbf{85.806}_{\pm 0.882}}$ $95.184_{\pm 0.216}$
Vicuna-13b $70.457_{\pm 2.444}$ $63.59_{\pm 2.804}$ $74.141_{\pm 2.241}$ $93.801_{\pm 0.507}$
Mixtral-Instruct-8x7b $80.862_{\pm 1.226}$ $70.697_{\pm 1.501}$ $85.725_{\pm 0.979}$ ${\textbf{95.686}_{\pm 0.232}}$
WizardLM-13b $74.923_{\pm 2.399}$ $66.468_{\pm 2.468}$ $77.355_{\pm 2.279}$ $92.759_{\pm 0.517}$
$\approx$70b Llama2-Chat-70b $89.567_{\pm 0.958}$ $83.374_{\pm 1.308}$ $90.416_{\pm 0.772}$ $96.436_{\pm 0.174}$
Qwen1.5-Chat-72b ${\textbf{99.154}_{\pm 0.348}}$ ${\textbf{98.412}_{\pm 0.54}}$ ${\textbf{99.138}_{\pm 0.286}}$ ${\textbf{99.757}_{\pm 0.072}}$
Platypus2-Instruct-70b $83.383_{\pm 2.235}$ $80.693_{\pm 2.39}$ $83.884_{\pm 2.125}$ $96.15_{\pm 0.463}$
Table 1: We scalably test the vulnerability of instruction-tuned LMs of different sizes against our attack. LMs with higher text similarity scores are more prone to output retrieved text verbatim. We show that LMs with stronger abilities are more vulnerable to prompt-injected data extraction: As model size increases, the maximum values for each size under each metric also increase. Notably, Llama2-Chat-7b can reach a ROUGE score over 80 and a BLEU score over 70.
5 Attacking Production LMs
In practice, users interact with more complex RAG systems, where the leakage problem can be mitigated by query filtering and output filtering. Besides, production LMs like ChatGPT are usually trained with alignment to detect such attacks and reject the query. For example, using regular prompt injection attack, we find that GPTs either output nothing or say “Sorry, I cannot fulfill that request”. In this section, we develop a prompt injection attack that can bypass the security guards of GPTs (OpenAI, 2023), a production RAG model built on ChatGPT, and successfully cause datastore leakage.
We start with a close inspection of GPTs’ system prompt (Appendix B.2). From the output, we notice that any customized GPT has at most four tools to use, among which myfiles_browser is used for browsing uploaded files, i.e. the customized datastore. Also, we find that GPTs use several function calls to do the retrieval work, one of which is search which “Runs a query over the file(s) uploaded in the current conversation and displays the results”. Therefore, we postulate that GPTs use an RIC-like method and explicitly ask GPTs to perform this function call by prompting with:
We refer readers to Appendix C for more examples of conducting the attack.
Experiment 1: Attack Success Rate.   We select 25 GPTs from the GPT store, spanning various data-sensitive domains including cyber security, law, finance, and medical. For each GPT, we generate the anchor queries by asking the GPT itself: “Generate some questions specific to your knowledge domain.” to simulate an adversary who has no prior knowledge of the datastore. After prompting all GPTs using the complete adversarial input, we report 100% attack success rate for datastore leakage on all the 25 GPTs, with 17 of them successfully attacked with 1 query and the rest succeeding with 2 queries. On average, we extract around 750 words from the datastore within each query.
Experiment 2: Reconstruction Rate.   We also investigate the possibility of reconstructing the entire customized datastore. We start with simulating a scenario where:
  • The datastore content might be included in the models’ pre-training data.
  • The attacker has partial prior knowledge about the datastore and thus can generate relevant queries.
We select a customized GPT built upon Harry Potter, [^1] https://chat.openai.com/g/g-TuM1IkwuA-harry-potter and its leaked system prompt shows that it uses the entire series of Harry Potter (7 books). Since the GPT outputs retrieved chunks in order, our adversary’s goal is to reconstruct the first book, Harry Potter and the Sorcerer’s Stone (77,000 words and 334,700 characters), by collecting the foremost output. An example of GPT output can be seen in Figure 4 in Appendix. To make anchor queries span a wide range of the book, we prompt the GPT with: “Generate 100 questions that cover each chapter of the book Harry Potter and the Sorcerer’s Stone”.
As a comparison, we simulate another more restricted yet realistic scenario with the following assumptions:
  • The datastore is constructed with knowledge that is not in the models’ pre-training data.
  • The attacker has no prior knowledge about the datastore and thus uses random queries for data extraction.
We make use of our collected latest Wikipedia corpus to build a new customized GPT. [^2] https://chat.openai.com/g/g-PorHEXuRq-wikigpt We generate anchor queries by prompting: “Generate 100 questions that cover most of your knowledge”. We then iteratively use each of the 100 questions as the anchor query to craft the model input and collect the output text. We found that for some queries, GPTs may retrieve overlapped text chunks. Removing duplicated chunks and concatenating all the chunks, we compute the reconstruction rate that measures how the extracted chunks reconstruct the original text data by calculating the ratio between the length of concatenation of deduplicated text chunks and that of the original text data.
Figure 2 shows that as we collect the GPT output with more queries, the reconstruction rate increases, and with only 100 questions in total, we can extract 41.73% text from the book and 3.22% text from our Wikipedia corpus. The reconstruction method could be potentially leveraged to audit a RAG system for copyrighted content. For example, copyright owners could craft diverse specific queries related to their works to reconstruct the datastore to check whether and how many of them have been included in the datastore.
Knowledge Size ROUGE-L BLEU F1 BERTScore
Wikipedia 7b $80.369_{\pm 1.679}$ $71.064_{\pm 2.033}$ $83.415_{\pm 1.375}$ $94.771_{\pm 0.301}$
13b $83.597_{\pm 1.104}$ $75.535_{\pm 1.404}$ $85.806_{\pm 0.882}$ $95.184_{\pm 0.216}$
70b $89.567_{\pm 0.958}$ $83.374_{\pm 1.308}$ $90.416_{\pm 0.772}$ $96.436_{\pm 0.174}$
Harry Potter 7b $92.815_{\pm 0.66}$ (+12.446) $81.818_{\pm 1.546}$ (+10.754) $90.023_{\pm 0.672}$ (+6.608) $95.581_{\pm 0.265}$ (+0.81)
13b $93.68_{\pm 0.805}$ (+10.083) $86.219_{\pm 1.374}$ (+10.684) $91.764_{\pm 0.834}$ (+5.958) $96.574_{\pm 0.213}$ (+1.39)
70b $95.31_{\pm 0.508}$ (+5.743) $88.276_{\pm 1.209}$ (+4.902) $92.897_{\pm 0.655}$ (+2.481) $96.957_{\pm 0.187}$ (+0.521)
Table 2: Ablation study on using different knowledge sources for Llama2-Chat models. We observe an apparent gain (Red) in text extraction for all 7b, 13b, and 70b models, leading us to hypothesize that LMs augmented with seen knowledge may be more prone to be extracted from the datastore.
Figure 2: Reconstruction rate of Harry Potter and the Sorcerer’s Stone (Blue) and Wikipedia (Green) against the number of domain-specific queries.
6 Ablation Studies
Instruction-tuning substantially enhances exploitability. We study how instruction tuning affects the vulnerability of data extraction (Figure 3). Still using our collected Wikipedia datastore, we compare the ROUGE score produced by the base model and the instruction-tuned model for Llama2-7b, Llama2-13b, Mistral-7b, and Mixtral-8x7b. On average, instruction tuning increases the ROUGE score between LM output under the attack and the retrieved context by 65.76. The large margins show that instruction tuning makes it easier to explicitly ask LMs to disclose their context, and this result aligns with our intuition that with strong instruction following ability, the LMs are also easier to be prompt injected, and thus malicious users can overwrite benign instructions and system prompts to cause unintended outputs.
Figure 3: Comparison of base and instruction-tuned LMs for Llama2-7b/13b, Mistral-7b, and Mixtral-8x7b.
Datastores are extractable if data are unseen during pre-training, and even more so if (likely) seen. Recall that we use the latest Wikipedia texts to make sure LMs have no prior knowledge about their datastore. As current models lack transparency in training data and contamination is widespread (Golchin and Surdeanu, 2023), it is unclear whether our results are an artifact of LMs’ pre-training data regurgitation, e.g. Harry Poter is likely already in the training data Books subset (Presser, 2020). We conduct experiments to control for such confounders and see how the knowledge source of the datastore would affect the data extraction of these open-sourced LMs. If an LM has seen the knowledge during the (pre-)training phase and we use the same knowledge as the datastore, we posit that it is more likely to generate such text verbatim. We choose Llama2-Chat as the model, use the original Harry Potter series as the knowledge source, and get anchor queries by asking GPT-4 to generate relevant questions. The results are shown in Table 2, with all other LMs’ settings remaining the same. On average, we observe gains of 9.42 for the ROUGE score, 8.78 for the BLEU score, 5.02 for the F1 score, and 0.91 for the BERTScore. Although we have no knowledge of Llama2’s training data, the gains in all four metrics shown above lead to a hypothesis that they have been trained on Harry Potter (possibly in the Books subset), which aligns with previous findings (Eldan and Russinovich, 2023; Reisner, 2024).
7 Conclusion
We investigate Prompt-Injected Data Extraction, an attack that extracts data from the datastore of a RAG system. Our study on both open-sourced and production RAG models reveals that instruction-tuned LMs are vulnerable to data extraction via copying their contexts, and we show that with stronger instruction-following capability, the vulnerability increases. We believe disclosing such problems can allow practitioners and policymakers aware of potential RAG safety and dual-use issues, and further contribute to the ongoing discussion on the regulation of generative models. Future work should incorporate different desiderata of multiple parties involved in emerging agent applications and RAG-enhanced production systems (Liu et al., 2023; Greshake et al., 2023) when diagnosing and mitigating data leakage of RAG datastore.
Acknowledgment
We thank Sizhe Chen, Robert Mahari, Rulin Shao for proofreading the draft. HZ is supported by an Eric and Susan Dunn Graduate Fellowship. SK acknowledges support from the Office of Naval Research under award N00014-22-1-2377 and the National Science Foundation Grant under award #IIS 2229881. This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence.
Appendix A Additional Experiment Details
A.1 Implementation
We use BM25 (Robertson et al., 2009) as the retriever. We use APIs provided by Together AI to perform inference and the hyperparameters we use for all instruction-tuned LMs are shown in Table 3 below.
Field Value
LLM Configurations
max_new_tokens 512
temperature 0.2
do_sample True
top_k 60
top_p 0.9
num_beams 1
repetition_penalty 1.8
Retriever Configurations
num_document 1
max_retrieval_seq_length 256
stride 128
Table 3: Default hyperparameters.
As for querying GPTs, we only use 100 questions to collect responses because the daily usage limit of GPTs is low. The Harry Potter GPT [^3] https://chat.openai.com/g/g-TuM1IkwuA-harry-potter and our WikiGPT [^4] https://chat.openai.com/g/g-PorHEXuRq-wikigpt are both available on the GPTs store. The ground truth text file we used to reconstruct Harry Potter GPT’s datastore is also publicly available. [^5] https://www.kaggle.com/datasets/moxxis/harry-potter-lstm
We use Huggingface’s evaluate module for computing ROUGE, BLEU, and BERTScore, and use NLTK’s ngrams and tokenize to compute token-level F1 score.
The 25 GPTs we successfully attack are categorized into 5 domains including finance, medical, etc, as shown in Table 4.
Domain Link
Cyber Security https://chat.openai.com/g/g-U5ZnmObzh-magicunprotect
https://chat.openai.com/g/g-b69I3zwKd-cyber-security-career-mentor
https://chat.openai.com/g/g-aaNx59p4q-hacktricksgpt
https://chat.openai.com/g/g-IZ6k3S4Zs-mitregpt
https://chat.openai.com/g/g-UKY6elM2U-zkgpt
https://chat.openai.com/g/g-HMwdSfFQS-secure-software-development-framework-ssdf-agent
https://chat.openai.com/g/g-qD3Gh3pxi-devsecops-guru
https://chat.openai.com/g/g-id7QFPVtw-owasp-llm-advisor
Law https://chat.openai.com/g/g-LIb0ywaxQ-u-s-immigration-assistant
https://chat.openai.com/g/g-w6KMGsg1K-bruno-especialista-en-lomloe
https://chat.openai.com/g/g-eDGmfjZb3-kirby
https://chat.openai.com/g/g-EznQie7Yv-u-s-tax-bot
https://chat.openai.com/g/g-0kXu7QuRD-leisequinha
https://chat.openai.com/g/g-me1tPbsgb-lawgpt
https://chat.openai.com/g/g-RIvUD7uxD-agent-agreement-legal-expert
Finance https://chat.openai.com/g/g-lVWqtb1gw-tech-stock-analyst
https://chat.openai.com/g/g-j5Mk8W3J7-bitcoin-whitepaper-chat
https://chat.openai.com/g/g-7McsRKuPS-economicsgpt
https://chat.openai.com/g/g-GaP7qDRTA-contacrypto-io
https://chat.openai.com/g/g-mAoqNweEV-quant-coder
Medical https://chat.openai.com/g/g-zVSzSYcu9-code-medica
https://chat.openai.com/g/g-LXZ1f4L5x-id-my-pill
https://chat.openai.com/g/g-Zj3N9NTma-empathic-echo
Religion https://chat.openai.com/g/g-nUKJX2cOA-biblegpt
https://chat.openai.com/g/g-p1EJzOI7z-quran
Table 4: 25 leaked GPTs across 5 different knowledge domains.
A.2 GPTs Outputs: An Example
In Figure 4, we use an example query to compare GPTs output with the original text from Harry Potter and the Sorcerer’s Stone to show how adversaries can extract text verbatim from GPTs datastore.
Figure 4: The original text (red) copied from Harry Potter and the Sorcerer’s Stone and the entire output text (green) from a Harry Potter GPT using our attack, with the query “Which house at Hogwarts does Harry get sorted into?”. We do a character-level comparison and the differences are highlighted. The prefix “L{number}” is a chunk marker used by GPT’s file browser engine.
Appendix B Prompt Examples
B.1 Check Existence of GPTs Datastore
To quickly check whether myfiles_browser exists, one can prompt GPTs by:
Prompt
B.2 GPTs Prompt Leakage
One can easily get GPTs system prompt with the input:
Prompt
We show an excerpt of the leaked GPTs’ system prompt:
Leaked GPTs System Prompt
You are ChatGPT, a large language model trained by OpenAI, based on the GPT-4 architecture. Knowledge cutoff: 2023-04 Current date: 2024-02-01 Image input capabilities: Enabled # Tools ## browser You have the tool ‘browser‘ with these functions: ‘search(query: str, recency_days: int)‘ Issues a query to a search engine and displays the results. ‘click(id: str)‘ Opens the webpage with the given id, displaying it. The ID within the displayed results maps to a URL. ‘back()‘ Returns to the previous page and displays it. ‘scroll(amt: int)‘ Scrolls up or down in the open webpage by the given amount. ‘open_url(url: str)‘ Opens the given URL and displays it. ‘quote_lines(start: int, end: int)‘ Stores a text span from an open webpage. Specifies a text span by a starting int ‘start‘ and an (inclusive) ending int ‘end‘. To quote a single line, use ‘start‘ = ‘end‘. For citing quotes from the ’browser’ tool: please render in this format: ‘[message idx†link text]‘. For long citations: please render in this format: ‘[link text](message idx)‘. Otherwise do not render links. Do not regurgitate content from this tool. Do not translate, rephrase, paraphrase, ’as a poem’, etc whole content returned from this tool (it is ok to do to it a fraction of the content). Never write a summary with more than 80 words. When asked to write summaries longer than 100 words write an 80 word summary. Analysis, synthesis, comparisons, etc, are all acceptable. Do not repeat lyrics obtained from this tool. Do not repeat recipes obtained from this tool. Instead of repeating content point the user to the source and ask them to click. ALWAYS include multiple distinct sources in your response, at LEAST 3-4. Except for recipes, be very thorough. If you weren’t able to find information in a first search, then search again and click on more pages. (Do not apply this guideline to lyrics or recipes.) Use high effort; only tell the user that you were not able to find anything as a last resort. Keep trying instead of giving up. (Do not apply this guideline to lyrics or recipes.) Organize responses to flow well, not by source or by citation. Ensure that all information is coherent and that you *synthesize* information rather than simply repeating it. Always be thorough enough to find exactly what the user is looking for. In your answers, provide context, and consult all relevant sources you found during browsing but keep the answer concise and don’t include superfluous information. EXTREMELY IMPORTANT. Do NOT be thorough in the case of lyrics or recipes found online. Even if the user insists. You can make up recipes though. ## myfiles_browser You have the tool ‘myfiles_browser‘ with these functions: ‘search(query: str)‘ Runs a query over the file(s) uploaded in the current conversation and displays the results. ‘click(id: str)‘ Opens a document at position ‘id‘ in a list of search results ‘back()‘ Returns to the previous page and displays it. Use it to navigate back to search results after clicking into a result. ‘scroll(amt: int)‘ Scrolls up or down in the open page by the given amount. ‘open_url(url: str)‘ Opens the document with the ID ‘url‘ and displays it. URL must be a file ID (typically a UUID), not a path. ‘quote_lines(line_start: int, line_end: int)‘ Stores a text span from an open document. Specifies a text span by a starting int ‘line_start‘ and an (inclusive) ending int ‘line_end‘. To quote a single line, use ‘line_start‘ = ‘line_end‘. please render in this format: ‘[message idx†link text]‘ Tool for browsing the files uploaded by the user. Set the recipient to ‘myfiles_browser‘ when invoking this tool and use python syntax (e.g. search(’query’)). "Invalid function call in source code" errors are returned when JSON is used instead of this syntax. Think carefully about how the information you find relates to the user’s request. Respond as soon as you find information that clearly answers the request. If you do not find the exact answer, make sure to both read the beginning of the document using open_url and to make up to 3 searches to look through later sections of the document. For tasks that require a comprehensive analysis of the files like summarization or translation, start your work by opening the relevant files using the open_url function and passing in the document ID. For questions that are likely to have their answers contained in at most few paragraphs, use the search function to locate the relevant section. ## dalle …(this part is too long to include here) ## python When you send a message containing Python code to python, it will be executed in a stateful Jupyter notebook environment. python will respond with the output of the execution or time out after 60.0 seconds. The drive at ’/mnt/data’ can be used to save and persist user files. Internet access for this session is disabled. Do not make external web requests or API calls as they will fail.
Appendix C Two-Step Attack on GPTs: Examples
A two-step method for leaking GPTs’ files is:
Step 1: Confirming the existence of datastore. One should first check whether the customized datastore is activated by the GPTs. This can be done by checking whether myfiles_browser is shown under the heading Tools in the leaked system prompt. We refer readers to Appendix B.1 for more information on how to craft the prompt.
Step 2: Attacking the datastore. Simply prompt GPTs using the attack above with an anchor query. If an adversary has no prior knowledge about the GPTs, anchor queries can be generated by asking the GPTs to randomly generate questions relevant to their knowledge.
As an example, the “Harry potter” GPT can be attacked as follows (Figure 5):
Figure 5: An excerpt of a conversation with Harry potter GPT.
In Figure 6, we take another customized GPT called “EconomicsGPT” as an example to show how to elicit the datastore content. [^6] https://chat.openai.com/g/g-7McsRKuPS-economicsgpt In Figure 7, we use a GPT called “zkGPT” to show how we can perform the extraction attack even when we don’t have prior knowledge about the datastore. [^7] https://chat.openai.com/g/g-UKY6elM2U-zkgpt
Figure 6: An excerpt of a conversation with EconomicsGPT.
Figure 7: An excerpt of a conversation with zkGPT.
Note that the output format varies: sometimes GPTs use json and sometimes output text as chunks as shown here. In some cases, one might need to ask the GPT to regenerate due to “No results found” related output or modify the anchor query.
Also, sometimes GPTs cannot find relevant results. One can try modifying the anchor query by making it longer and richer in information.